OpenStudy (anonymous):

find cosθ if sinθ=-4/5 and is in quad three

OpenStudy (solomonzelman):

can you find the third side?

OpenStudy (anonymous):

hey it doesnt matter but from what i seen in the past.. try to construct that triangle using the reference angle. @SolomonZelman

OpenStudy (solomonzelman):

will see, this time it is not a necessity though, I see you agree to this. it sounds like you are criticizing me, jk but I know I suck at math....

OpenStudy (solomonzelman):

there is a rule: \(\large\color{black}{ a^2+b^2=c^2 }\) (where c is the hypotenuse) plugging your values: \(\large\color{black}{ a^2+4^2=5^2 }\) fiond te 3rd side.

OpenStudy (anonymous):

well with that many fans and great testimonials i say you rock!!!!! :D... but anyway the reason i suggested it is because im pretty sure he/she is at that stage where their learning co-terminal and reference angles.

OpenStudy (anonymous):

your right @magepker728

OpenStudy (anonymous):

but im still lost

OpenStudy (anonymous):

ok then sinetheta is opposite/hypotenuse in this case u have sinetheta in 3rd quadrant

OpenStudy (anonymous):

and costheta is adjecent/hypotenuse we already know the hypotenuse so adjacent is missing .. to find it used the equations salomonzelmen provided

OpenStudy (anonymous):

so in other words find a